OpenStudy (anonymous):

Solve for u: 3u^2+5=-16u If there is more than one solution separate them with commas.

OpenStudy (agreene):

\[3 u^2+5 = -16 u\] Subtract -16 u from both sides: \[3 u^2+16 u+5 = 0\] Divide both sides by 3: \[u^2+\frac{16 u}{3}+\frac{5}{3} = 0\] Subtract 5/3 from both sides: \[u^2+\frac{16 u}{3} = -\frac{5}{3}\] Add 64/9 to both sides: \[u^2+\frac{16 u}{3}+\frac{64}{9} = \frac{49}{9}\] Factor the left hand side: \[(u+\frac{8}{3})^2 = \frac{49}{9}\] Take the square root of both sides: \[|u+\frac{8}{3}| = \frac{7}{3}\] Eliminate the absolute value: u+8/3 = -7/3 or u+8/3 = 7/3 Subtract 8/3 from both sides: u = -5 or u+8/3 = 7/3 Subtract 8/3 from both sides: \[\huge u = -5, \frac{1}{3}\]